# AI-based Drone Assisted Human Rescue in Disaster Environments: Challenges and Opportunities

### v1

@@ I Introduction
Related Work: Several works have been undertaken in the field of search and rescue operations. Alsamhi et al. [alsamhi2022uav], assess the network performance of UAV-supported intelligent edge computing to accelerate Search And Rescue (SAR) operations. This technology offers rapid deployment capabilities and can aid in disaster-related rescues.
Alawad et al. [alawad2023unmanned], propose a crisis and disaster management system that leverages a swarm optimization algorithm (SOA) to enhance disaster and crisis management efforts. Within this system, the UAV search and rescue team operates under the delay tolerant network (DTN) strategy, enabling efficient exploration capabilities.
Erdelj and Natalizio [erdelj2016uav], identify the primary applications of UAV networks in disaster management and examine key research challenges within the domain of UAV-assisted disaster management.
In our previous work, [kulhandjian2023ai], we have created an integrated system featuring two distinct sensing mechanisms, enabling real-time detection and precise localization of humans and animals even in thick smoke, enhancing the situational awareness of firefighters on-site. An extended version of this work has been submitted for a patent [kulhandjian2022system].
In this article, we illustrate our strategy for combatting large-scale wildfires through the utilization of unmanned aerial vehicles (UAVs) and neural network algorithms, with a primary focus on significantly reducing human casualties [Albanese2021]. UAV technology plays a pivotal role in mitigating the impact of emergency situations and expediting rescue and recovery efforts [Namuduri2017, Tomic2012, Grover2017]. Our foremost objective is the swift detection of emerging disasters and prompt reporting to all relevant law enforcement agencies. A swift response to such incidents presents a valuable opportunity to contain the wildfire’s spread, with every moment gained being crucial in limiting the disaster’s extent.
@@ VI-C Audio/Microphone Array
While the majority of research efforts have centered on the development of video-based solutions, UAV-embedded audio-based localization has garnered comparatively less attention [Basiri2012, Strauss2018]. Nevertheless, the utilization of UAVs equipped with microphone arrays holds the potential to be of paramount importance in locating individuals during emergency situations. This potential becomes particularly evident when video sensors are hindered by factors such as poor lighting conditions (e.g., at night or in fog) or obstacles that restrict their field of view [Deleforge2019].
@@ VI-D Radars
Intelligent onboard sensors, such as cameras, microphone arrays, and radar, which are particularly effective during the early stages of disaster response. However, radar’s unique ability to penetrate through objects and operate in low-visibility conditions makes it indispensable for detecting occluded human subjects, both on and under debris, when other sensing abilities may falter [Rong2021]. A radar is an electromagnetic sensor utilized for the detection, localization, tracking, and identification of various objects over significant distances. Its operation involves the transmission of electromagnetic energy towards objects, commonly referred to as targets, and the subsequent observation of the echoes reflected from these objects. Radar not only determines the presence, location, and speed of these objects but in certain cases can also ascertain their size and shape. What sets radar apart from optical and infrared sensors is its ability to detect distant objects under adverse weather conditions and accurately measure their distance.
@@ VIII-A Synchronization Among UAVs
UAVs, in general, offer advantages such as rapid deployment, flexible reconfiguration, and improved communication capabilities due to their short-range line-of-sight links. Nevertheless, the deployment of highly mobile and energy-constrained UAVs for wireless communications also brings forth a host of new challenges [Zeng2016].
UAVs have become indispensable in modern life, and achieving precise time synchronization is crucial for multi-UAV flight. However, the current approach, which relies on GNSS satellite navigation systems or ground control stations for time synchronization, has limitations and poses significant security risks [Panda2019]. The challenge lies in achieving highly accurate time synchronization for all drones within the network independently, without relying on external time sources.
In the context of a distributed architecture, such as a large-scale UAV formation, the traditional master-slave time synchronization method is no longer suitable. Instead, the firefly synchronization model, which has ancient origins and has been studied in various fields like biology, chemistry, and mathematics, offers a novel approach to address the issue of distributed time synchronization. The concept involves each UAV transmitting its current time information [Liu2015]. Once neighboring nodes receive this information, they perform a straightforward arithmetic averaging operation. This calculated average is then used as the countdown for the next transmission. This process is iterated several times until all nodes in the network eventually converge to identical clock values on average. This achieves distributed time synchronization for the entire formation network as depicted in Fig. 9.
@@
In a UAV formation, multiple UAVs are interconnected and can interact with one another [Alsolami2021]. Each UAV serves as a node within the network, establishing a decentralized network structure without a central node. The relative positioning between these nodes is consistent.
@@ VIII-B UAV Network Security
In recent years, there has been a notable increase in the malicious use of unmanned aerial vehicles (UAVs). These attacks have become more frequent and can have severe consequences. As a result, various industries and standardization bodies are actively exploring ways to enhance the security of UAV systems and networks [KaiYun2022].
To address this issue comprehensively, threats and protective measures are categorized based on the characteristics of the first four layers of the OSI model: the physical layer, data link layer, network layer, and transport layer. In order to provide a deeper insight, the security mechanisms under examination are thoroughly assessed in terms of their security requirements and objectives [Erdelj2017]. These objectives encompass aspects such as availability, authentication, authorization, confidentiality, integrity, and non-repudiation.
@@ IX Design Architecture
Numerous advanced technology applications, such as employing a hexacopter for various purposes, necessitate a well-structured architecture comprising both hardware and software modules [Hu2021]. In the case of a hexacopter, this layered organization is instrumental in facilitating efficient functionality and task coordination. The design encompasses various hardware and software components that collaboratively operate to accomplish predefined objectives.
@@ IX-C Handling UAV Failures
While there exists extensive literature on fault tolerance methods applied to UAVs [Erdos2013], the field of failure prevention techniques is relatively less detailed and somewhat outdated. Various methods, including fault tree analysis (FTA), failure modes and effects analysis (FMEA), and component analysis (CA), are employed in this context. FTA serves as a qualitative assessment method to guide designers, planners, or operators in understanding potential system failures and devising strategies to address their causes. It typically complements FMEA, which involves an exhaustive examination of failure modes for each system component, assessing their impact on system operation. FMEA is most effective when implemented during the design phase, allowing critical failure probabilities to be identified and mitigated early in the process.
@@ XI Conclusion
[…]
[Uncaptioned image] Michel Kulhandjian (M’18-SM’20) received his B.S. degree in Electronics Engineering and Computer Science (Minor), with “Summa Cum Laude” from the American University in Cairo (AUC) in 2005, and the M.S. and Ph.D. degrees in Electrical Engineer from the State University of New York at Buffalo in 2007 and 2012, respectively. He was employed at Alcatel-Lucent, in Ottawa, Ontario, in 2012. In the same year he was appointed as a Research Associate at EION Inc. In 2016 he was appointed as a Research Scientist at the School of Electrical Engineering and Computer Science at the University of Ottawa. He was also employed as a senior embedded software engineer at L3Harris Technologies from 2016 to 2021. Currently, he is a Research Scientist at the Electrical and Computer Engineering Department at Rice University. He received the Natural Science and Engineering Research Council of Canada (NSERC) Industrial R&D Fellowship (IRDF). His research interests include wireless multiple access communications, adaptive coded modulation, waveform design for overloaded code-division multiplexing applications, RF and audio fingerprinting, channel coding, space-time coding, adaptive multiuser detection, statistical signal processing, machine learning, covert communications, spread-spectrum steganography, and steganalysis. He actively serves as a member of the Technical Program Committee (TPC) of IEEE WCNC, IEEE GLOBECOM, IEEE ICC, and IEEE VTC, among others. In addition, he serves as an Associate Editor of Annals of Telecommunications Journal, a guest editor for the Journal of Sensor and Actuator Networks (JSAN), and a Review Editor at Frontiers in Communications and Networks. He is a recipient of the best paper award at the 48th Wireless World Research Forum (WWRF) in 2022.
[Uncaptioned image] Hovannes Kulhandjian (S’14-M’15-SM’20) received the B.S. degree (magna cum laude) in electronics engineering from The American University in Cairo, Cairo, Egypt, in 2008, and the M.S. and Ph.D. degrees in electrical engineering from the State University of New York at Buffalo, Buffalo, NY, USA, in 2010 and 2014, respectively. From December 2014 to July 2015, he was an Associate Research Engineer with the Department of Electrical and Computer Engineering, Northeastern University, Boston, MA, USA. He is currently an Associate Professor with the Department of Electrical and Computer Engineering, California State University, Fresno, Fresno, CA, USA. His current research interests include wireless communications and networking, with applications to underwater acoustic communications, visible light communications, and applied machine learning. He has served as a guest editor for IEEE Access - Special Section Journal on Underwater Wireless Communications and Networking. He has also served as a Session Co-Chair for IEEE UComms 2020 and session Chair for ACM WUWNet 2019. He actively serves as a member of the Technical Program Committee for ACM and IEEE conferences such as IEEE GLOBECOM 2015-2022, UComms 2022, PIMRC 2020, WD 2019, ACM WUWNet 2019, ICC 2015-2023, among others.
[Uncaptioned image] Levon Hakob Aslanyan received the B.S. degree from the Novosibirsk State University in 1968, and the Ph.D. and Doctor of Science (Russian) degrees from the Novosibirsk State University, in 1976 and 1997, respectively. He has been promoted to Professor position since 1997. He has been the corresponding member of the National Academy of Sciences of the Republic of Armenia (NAS RA), since 2014. Currently, he heads the Department of Discrete Modeling, Analysis, and Recognition Technologies at the Institute for Informatics and Automation Problems of the NAS RA. His research interests are in mathematical logic, discrete mathematics, the mathematical theory of pattern recognition, and artificial intelligence. He has made significant contributions in the field of complexity studies of loosely defined Boolean functions, describing sets encompassing all solutions to discrete isoperimetry and discrete tomography problems, advancing logic-combinatorial pattern recognition theory, and contributing to the field of bioinformatics. He authored more than $200200$ papers and supervised $22$ doctoral and $1414$ Ph.D. students.

### v2

@@ I Introduction
Related Work: Several works have been undertaken in the field of search and rescue operations. Alsamhi et al. [1], assess the network performance of UAV-supported intelligent edge computing to accelerate Search And Rescue (SAR) operations. This technology offers rapid deployment capabilities and can aid in disaster-related rescues.
Alawad et al. [2], propose a crisis and disaster management system that leverages a swarm optimization algorithm (SOA) to enhance disaster and crisis management efforts. Within this system, the UAV search and rescue team operates under the delay tolerant network (DTN) strategy, enabling efficient exploration capabilities.
Erdelj and Natalizio [3], identify the primary applications of UAV networks in disaster management and examine key research challenges within the domain of UAV-assisted disaster management.
In our previous work, [4], we have created an integrated system featuring two distinct sensing mechanisms, enabling real-time detection and precise localization of humans and animals even in thick smoke, enhancing the situational awareness of firefighters on-site. An extended version of this work has been submitted for a patent [5].
In this article, we illustrate our strategy for combatting large-scale wildfires through the utilization of unmanned aerial vehicles (UAVs) and neural network algorithms, with a primary focus on significantly reducing human casualties [6]. UAV technology plays a pivotal role in mitigating the impact of emergency situations and expediting rescue and recovery efforts [7, 8, 9]. Our foremost objective is the swift detection of emerging disasters and prompt reporting to all relevant law enforcement agencies. A swift response to such incidents presents a valuable opportunity to contain the wildfire’s spread, with every moment gained being crucial in limiting the disaster’s extent.
@@ VI-C Audio/Microphone Array
While the majority of research efforts have centered on the development of video-based solutions, UAV-embedded audio-based localization has garnered comparatively less attention [10, 11]. Nevertheless, the utilization of UAVs equipped with microphone arrays holds the potential to be of paramount importance in locating individuals during emergency situations. This potential becomes particularly evident when video sensors are hindered by factors such as poor lighting conditions (e.g., at night or in fog) or obstacles that restrict their field of view [12].
@@ VI-D Radars
Intelligent onboard sensors, such as cameras, microphone arrays, and radar, which are particularly effective during the early stages of disaster response. However, radar’s unique ability to penetrate through objects and operate in low-visibility conditions makes it indispensable for detecting occluded human subjects, both on and under debris, when other sensing abilities may falter [13]. A radar is an electromagnetic sensor utilized for the detection, localization, tracking, and identification of various objects over significant distances. Its operation involves the transmission of electromagnetic energy towards objects, commonly referred to as targets, and the subsequent observation of the echoes reflected from these objects. Radar not only determines the presence, location, and speed of these objects but in certain cases can also ascertain their size and shape. What sets radar apart from optical and infrared sensors is its ability to detect distant objects under adverse weather conditions and accurately measure their distance.
@@ VIII-A Synchronization Among UAVs
UAVs, in general, offer advantages such as rapid deployment, flexible reconfiguration, and improved communication capabilities due to their short-range line-of-sight links. Nevertheless, the deployment of highly mobile and energy-constrained UAVs for wireless communications also brings forth a host of new challenges [14].
UAVs have become indispensable in modern life, and achieving precise time synchronization is crucial for multi-UAV flight. However, the current approach, which relies on GNSS satellite navigation systems or ground control stations for time synchronization, has limitations and poses significant security risks [15]. The challenge lies in achieving highly accurate time synchronization for all drones within the network independently, without relying on external time sources.
In the context of a distributed architecture, such as a large-scale UAV formation, the traditional master-slave time synchronization method is no longer suitable. Instead, the firefly synchronization model, which has ancient origins and has been studied in various fields like biology, chemistry, and mathematics, offers a novel approach to address the issue of distributed time synchronization. The concept involves each UAV transmitting its current time information [16]. Once neighboring nodes receive this information, they perform a straightforward arithmetic averaging operation. This calculated average is then used as the countdown for the next transmission. This process is iterated several times until all nodes in the network eventually converge to identical clock values on average. This achieves distributed time synchronization for the entire formation network as depicted in Fig. 9.
@@
In a UAV formation, multiple UAVs are interconnected and can interact with one another [17]. Each UAV serves as a node within the network, establishing a decentralized network structure without a central node. The relative positioning between these nodes is consistent.
@@ VIII-B UAV Network Security
In recent years, there has been a notable increase in the malicious use of unmanned aerial vehicles (UAVs). These attacks have become more frequent and can have severe consequences. As a result, various industries and standardization bodies are actively exploring ways to enhance the security of UAV systems and networks [18].
To address this issue comprehensively, threats and protective measures are categorized based on the characteristics of the first four layers of the OSI model: the physical layer, data link layer, network layer, and transport layer. In order to provide a deeper insight, the security mechanisms under examination are thoroughly assessed in terms of their security requirements and objectives [19]. These objectives encompass aspects such as availability, authentication, authorization, confidentiality, integrity, and non-repudiation.
@@ IX Design Architecture
Numerous advanced technology applications, such as employing a hexacopter for various purposes, necessitate a well-structured architecture comprising both hardware and software modules [20]. In the case of a hexacopter, this layered organization is instrumental in facilitating efficient functionality and task coordination. The design encompasses various hardware and software components that collaboratively operate to accomplish predefined objectives.
@@ IX-C Handling UAV Failures
While there exists extensive literature on fault tolerance methods applied to UAVs [21], the field of failure prevention techniques is relatively less detailed and somewhat outdated. Various methods, including fault tree analysis (FTA), failure modes and effects analysis (FMEA), and component analysis (CA), are employed in this context. FTA serves as a qualitative assessment method to guide designers, planners, or operators in understanding potential system failures and devising strategies to address their causes. It typically complements FMEA, which involves an exhaustive examination of failure modes for each system component, assessing their impact on system operation. FMEA is most effective when implemented during the design phase, allowing critical failure probabilities to be identified and mitigated early in the process.